OpenStudy (anonymous):

Sams calculator represented his answer as 6.987 E-04. Rewrite his answer in standard form.

Directrix (directrix):

6.987 * 10^( -4)= = .0006987

OpenStudy (anonymous):

what does ^ mean?

OpenStudy (anonymous):

please explain

Directrix (directrix):

^ is a carat symbol. In this context, it means raised to a power. Hence, 2^3 means 2 to the third power or 2³ = 2*2*2 = 8. 10 ^ (-4) = .0001. 1/10 * 1/10 * 1/10 * 1/10

OpenStudy (anonymous):

what does the e mean though

Directrix (directrix):

Exponent as in an exponent of -4.

OpenStudy (s):

e on the calculator means "times ten to the..." so 6.987 e -04 is the same as 6.987 * 10^-4 or 6.987 time ten to the power of negative four